Field and CPL Updates

2006 Greenland

July 2, 2006
Update provided by Jay Johnson (ICDS) and Geoff Hargreaves (NICL)
project update image

Activity: DISC Drill Test - Summit Greenland

Period: 6/26-7/01/06

Weather Conditions: Temperature range -25°:C to -5°:C, wind 3-21 knots

ICDS Personnel on Site:
Jay Johnson
Kristina Dahnert
Michael Jayred
Bill Mason
Nicolai Mortensen
Paul Sendelbach
Scott Hamon
John Fowler

Activities During Period:

Drilling - This week we drilled a total of 79.869 meters. The bore hole depth is 698.723 meters. Average core length was 2.41 meters. The longest core drilled this week was 2.567 meters. Between the depths of 638.35 and 633.72 three ~2 meter cores were saved by NICL to be possibly used for our brittle ice acceptance test. Our intent is two drill three more cores to replace these when we hit 730 meters. The first three cores were saved in case something was to happen causing us to discontinue drilling before reaching 730 meters. We are starting to see more cores with a fracture in them when they are pushed from the core barrel. However it is rare to see more than one fracture in a 2.5 meter core. We have yet to see core with poker chipping.

Cutters - I modified a set of cutters by adding a ~1mm radius to the inside corner. NICL reported seeing very little change in the number or orientation of surface fractures on the cores drilled with these cutters as compared to cores drilled with sharp corner cutters.

Core orientation - This week we tested and verified that we are able to orientate core breaks using instrumentation.

Instrument section J - We have been running this section all week. One of the micro controllers continues to require several attempts to come on line. We only have to power it up in the morning and then it's left on all day. Once it is booted up it will stay running. This instrument section has been working well otherwise.

Instrument section K - Nicolai spent most of the week inspecting, cleaning, and repairing the electronics. He has completed the repairs and it is ready to be tested. To do a full test we will have to install it in the drill.

Work Schedule - We switched back to one shift this week. This has put our drilling pace and fluid usage on track to complete the last of the drilling and testing by July 13th.

Power Monitoring - On Friday Nicolai set up a data logger in Lab View to log power usage from our CT monitor. Next week we will collect data.


Activity: NICL Greenland Test

Period: 6/26-7/02/06

NICL Personnel on Site: Geogg Hargreaves, Brian Bencivengo

Activities During Period:

6/26/06 Monday - Brian tested cutting blades and drop rates. Have found that the linear actuator is a benefit, it allows controlled drop rates, also frees up operator to do other things, or at least not have to stand and hold a bloody cold saw while it does it's work. This also allows us to test saw blade drop rates at a specified speed (no human error due to randomness of saw drop.)

6/27/06 Tuesday- Brian continues testing different saw blades and drop rates. Geoff tries to get printer/wireless working. Conducted Brittle Ice Acceptance Test (ICDS wanted to try one now in case the drill failed at some point. We will conduct another test when the depth is at least 730 meters.) ICDS passed this test with ease. Tested "vacuum wand" on smooth and broken ice (some that we had broken,) today. The wand worked quite well, and did not pick up any major ice pieces. The ice appeared to have more ridges (~1 mm deep,) on it today, but the microcracks along the core were not as bad.

6/28/06 Wednesday - Basement for core drying tests started. Brian helped with the digging/setup. Brian bags and stores core so we will have several cores to test in the drying booth. Geoff finally got the printer to go using USB connection to server, and printed some core cards. Tried "containing" ice that we broke with panty hose. It does not work so well (rips,) possibly due to the fact that the panty hose is for "svelte sized" women. We took a piece of core from after the FED as a control sample. ICDS switched cutter teeth to a smaller radius today. The microcracks appear to occur with increasing frequency, and there is bad pitting. Started wiping the core with towels after the FED to try and gauge wetness. The towels removed some Isopar, but it is still very hard to quantify the amount.

6/29/06 Thursday - Basement is finished, many thanks to the VECO folks who did the work to make it happen. We started placing core into this new storage area. Changed out suction inlet to R2D2 so that we can test that vacuum with the FED. Jay from ICDS turned the "barrel to FED bridge" down a bit so that the barrel end will fit into the barrel. This has proved a benefit, when trying to pass broken ice through the FED. Brian took a "wet" ice sample by cracking the top end of the core (cracking the first 20 cm or so of core,) which passed easily through the FED by using the Barrel FED bridge. Mounted the digital measuring stick, and began testing for long term temperature tolerance and calibration. ICDS tried to do an azimuth alignment with two cores. The azimuth was out of line by 180 degrees. We did another later in the day, and the azimuth was 90 degrees out of phase. We tested NICL's small VOC detectors, but they did not work with the Isopar (the detectors need the Isopar to be more volatile.) We looked into firn cooling the tent, but decide that it was not very feasible. The snow is very packed around the NICL tent (due to traffic, etc.) and there are power lines buried all around. ICDS tried moving around cutter heads to try and find the reason that every 4th cut is deeper than the other three. The first ~30-40 cm of the cores today also had a very deep (~2mm,) rough cut pattern. This is possibly due to the drill "pushing" into the core initially.

6/30/06 Friday - Geoff worked on emails, WAIS 90% drawings, and tried to cobble up a quickie database to test the usability of the digital measuring stick. Brian continued logging core, and took samples for 2, 4, 8, and 12 hour Isopar drying tests. All of the cores mated perfectly with the others today. Jay eased cutters towards the end of today and the rough ~30 cm is gone. The micro cracks and deeper 4th cut is still present, but this core is much nicer. After this, the core looked the best that it has in a week.

7/01/06 Saturday - Worked on unpacking and setting up shelving for drying booth. Today is the Summit July 4th celebration day, so knocked off early. Took tour of GRIP site main building. Worked on weekly report and entered data into computers tonight.

Comments (Problems, Concerns, Recommendations, etc.):

  • NICL could possibly measure temperatures faster with an infrared "temperature gun". As well, having thermometers on the Aluminum tray would be a great way to measure ice/tray temperatures. We are going to borrow one from the Mechanic Robin here at Summit and give it a try.
  • Opening the tent door seems to cool our tent down ~10 degrees. We can get about a ~12 degree cool down if ICDS runs their fans to draw cool air through our tent. We have also found that on days with a southwest wind, our tent fills with drift, and makes it dangerous (slippery) to work.
  • Both NICL and ICDS are finding that certain materials are intolerant of the Isopar. For instance, NICL has to replace the flange gasket on Oscar's knock-out-drum.
  • The "vacuum wand" needs to be just a shade wider, and some sort of flange would benefit it.
  • If we can mount a laser that projects a vertical line onto the Balluff measuring stick slider, we can project a beam onto the core making it easier to procure measurements as well as label the core.
  • NICL would like to have some way to know whether or not the drill is going up or down, and ETA of core when at WAIS.
  • Discussed with ICDS ways to receive data from their database. We will have to work with Paul and Nickoli to get this to work. It is not possible at this time. Neither of our databases are working as well as we would like.
  • SAW:
    • Spalling/chipping seen at the top of the core when cutting may be due to the fact that the saw blade is vibrating perpendicular to the core upon contact.
    • We have 0.5" width blades to test, but in order to put them onto the large saw (which is setup for 3/4" blades) we have to reconfigure it. We have decided to wait until after the next brittle ice test to do this or possible wait until return to Denver.
    • Tooth ratios of 6, 10, and 6-10 teeth (variable tooth) per inch seem to cut the ice very well at several different drop speeds. The blade speed we are running is about 970 SFPM.
  • The melt/slush buildup we were seeing earlier has disappeared, We think due primarialy to outside temps being cooler and keeping the door open with ICDS fans on we can keep our weather port cooler.
  • Core receiving system:
    • Placing the Aluminum and wood trays in the snow only cooled them ~1 degree (with tent cooled down.)
    • Need to place a thin slippery surface between the Aluminum tray and the 4 meter push out table. Sliding Aluminum on Aluminum is causing it to gall.
    • The FED, Aluminum tray, saw, roller track, and wood trays are having alignment problems. This is due to an uneven floor, and material warp. It appears that the FED may be losing its alignment due to its mounting method.
    • ICDS has had some problems pushing the core onto the Alumium V-tray. This could be due to FED misalignment or the core catching on the lay flat. We need to look into an interface of some type to make a smooth transition. Possibly a block of plastic or aluminum shaped to make the transition.
    • After conversation with Bill and Jay, we have decided to change the mounting method for the receiving table and Fed. The Table will be hard mounted to the Tower sections with 4 or 5 mounts, the FED will be mounted to the Tower sections with vertical and lateral adjustment to align the Fed to the 4m AL tray. The Barrel adjust will be the final alignment. This way there should be less play and less drift down the length of the 8m of barrel/FED/V tray unit.